OpenStudy (anonymous):

f(x)=(x^2-3x-18)/x+3

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

find the restriction on x factor numerator cancel factors graph the linear term with point of discontinuity at said restriction

OpenStudy (anonymous):

\[f(x)=\frac{ x^{2}-3x-18 }{ x+3 }\] From the start, in this form we can see that x can't equal -3. This would yield a solution of: \[x=-3\]\[f(-3) \rightarrow \frac{ 0 }{ 0 }\] Which is an indeterminate form. Factor the numerator and cancel: \[f(x)=\frac{ (x+3)(x-6) }{ x+3 }=x-6\] The graph is attached.

OpenStudy (anonymous):

sweeeeeeet. now I wonder if he got that